OpenStudy (anonymous):

Prove: tan^2 θ•cos^2 θ + cos^2 θ = 1. You must show all work.

OpenStudy (luigi0210):

Are those powers or angles?

OpenStudy (anonymous):

powers sorry

OpenStudy (xapproachesinfinity):

Double angles or power?

OpenStudy (luigi0210):

Use: \(\Large \frac{sin^2\theta}{cos^2\theta}=tan^2\theta\)

OpenStudy (anonymous):

How do I do that

OpenStudy (xapproachesinfinity):

let me draw something for you

OpenStudy (xapproachesinfinity):